# LLM-Assisted Static Analysis for Detecting Security Vulnerabilities

### v2

@@ Abstract
Software is prone to security vulnerabilities. Program analysis tools to detect them have limited effectiveness in practice due to their reliance on human labeled specifications. Large language models (or LLMs) have shown impressive code generation capabilities but they cannot do complex reasoning over code to detect such vulnerabilities especially since this task requires whole-repository analysis. We propose IRIS, a neuro-symbolic approach that systematically combines LLMs with static analysis to perform whole-repository reasoning for security vulnerability detection. Specifically, IRIS leverages LLMs to infer taint specifications and perform contextual analysis, alleviating needs for human specifications and inspection. For evaluation, we curate a new dataset, CWE-Bench-Java, comprising 120 manually validated security vulnerabilities in real-world Java projects. A state-of-the-art static analysis tool CodeQL detects only 27 of these vulnerabilities whereas IRIS with GPT-4 detects 55 ($+28$) and improves upon CodeQL’s average false discovery rate by 5% points. Furthermore, IRIS identifies 6 previously unknown vulnerabilities which cannot be found by existing tools.
@@ 1 Introduction
Security vulnerabilities pose a major threat to the safety of software applications and its users. In 2023 alone, more than 29,000 CVEs were reported—almost 4000 higher than in 2022 [9]. Detecting vulnerabilities is extremely challenging despite advances in techniques to uncover them. A promising such technique called static taint analysis is widely used in popular tools such as GitHub CodeQL [1], Facebook Infer [11], Checker Framework [4], and Snyk Code [39]. These tools, however, face several challenges that greatly limit their effectiveness and accessibility in practice.
@@
False positives due to lack of precise context-sensitive and intuitive reasoning. Second, it is well-known that static analysis often suffers from low precision, i.e., it may generate many false alarms [24, 21]. Such imprecision stems from multiple sources. For instance, the source or sink specifications may be spurious, or the analysis may over-approximate over branches in code or possible inputs. Further, even if the specifications are correct, the context in which the detected source or sink is used may not be exploitable. Hence, a developer may need to triage through several potentially false security alerts, wasting significant time and effort.
Limitations of prior data-driven approaches to improve static taint analysis. Many techniques have been proposed to address the challenges of static taint analysis. For instance, [34] proposed a probabilistic approach, MERLIN, to automatically mine taint specifications. A more recent work, Seldon [6], improves the scalability of this approach by formulating the taint specification inference problem as a linear optimization task. However, such approaches rely on analyzing the code of third-party libraries to extract specifications, which is expensive and hard to scale. Researchers have also developed statistical and learning-based techniques to mitigate false positive alerts [23, 18, 36]. However, such approaches still have limited effectiveness in practice [24].
Large Language Models (or LLMs) have made impressive strides in code generation and summarization. LLMs have also been applied to code related tasks such as program repair [45], code translation [35], test generation [27], and static analysis [28]. Recent studies [42, 25] evaluated LLMs’ effectiveness at detecting vulnerabilities at the method level and showed that LLMs fail to do complex reasoning with code, especially because it depends on the context in which the method is used in the project. On the other hand, recent benchmarks like SWE-Bench [20] show that LLMs are also poor at doing project-level reasoning. Hence, an intriguing question is whether LLMs can be combined with static analysis to improve their reasoning capabilities. In this work, we answer this question in the context of vulnerability detection and make the following contributions:
@@
Dataset. We curate a dataset of manually vetted and compilable Java projects, CWE-Bench-Java, containing 120 vulnerabilities (one per project) across four common vulnerability classes. The projects in the dataset are complex, containing 300K lines of code on average, and 10 projects with more than a million lines of code, making it a challenging benchmark for vulnerability detection. Our code and dataset are in the supplementary material and will be open-sourced upon publication.
Results. We evaluate IRIS on CWE-Bench-Java using 8 diverse open- and closed-source LLMs. Overall, IRIS obtains the best results with GPT-4, detecting 55 vulnerabilities, which is 28 ($103.7\%$) more than CodeQL, the existing best-performing static analyzer. We show that the increase is not at the expense of false positives, as IRIS with GPT-4 achieves an average false discovery rate of $84.82\%$, which is $5.21\%$ lower than that of CodeQL. Further, when applied to the latest versions of 30 Java projects, IRIS with GPT-4 discovered 6 previously unknown vulnerabilities.
@@ 2 Motivating Example
In contrast, IRIS takes a different approach by inferring project- and vulnerability-specific specifications on-the-fly by using LLMs. The LLM-based components in IRIS correctly identify the untrusted source and the vulnerable sink. IRIS augments CodeQL with these specifications and successfully detects the unsanitized data-flow path between the detected source and sink in the repository. However, augmented CodeQL produces many false positives, which are hard to eliminate using logical rules. To solve this challenge, IRIS encodes the detected code paths and the surrounding context into a simple prompt and uses an LLM to classify it as true or false positive. Specifically, out of 8 paths reported by static analysis, 5 false positives are filtered out, leaving the path in Fig. 2 as one of the final alarms. Overall, we observe that IRIS can detect many such vulnerabilities that are beyond the reach of CodeQL-like static analysis tools, while keeping false alarms to a minimum.
@@ 3.3 Inferring Taint Specifications using LLMs
Due to the sheer number of APIs to be labeled, we insert a batch of APIs in a single prompt and ask the LLM to respond with JSON formatted strings. The batch size is a tunable hyper-parameter. We adopt few-shot (usually 3-shot) prompting strategy for labeling external APIs, while zero-shot is used for labeling internal APIs. Notably for internal APIs, we also include information from repository readme and JavaDoc documentations, if applicable. In practice, we find that this extra information helps LLM understand the high-level purpose and usage of the codebase, resulting in better labeling accuracy. Due to space limitation, we leave the full prompt templates and other implementations details in the Appendix. At the end of this stage, we have successfully obtained $\bm{S}_{\textit{source}}^{C}$ and $\bm{S}_{\textit{sink}}^{C}$ which are going to be used by the static analysis engine in the next stage.
@@ 3.4 Vulnerability Detection
$\textit{CodeQL}(\mathbb{G}^{P},\bm{S}_{\textit{source}}^{C},\bm{S}_{\textit{sink}}^{C},\textit{Query}^{C})=\{\textit{Path}_{1},\ldots,\textit{Path}_{k}\}$
CodeQL itself contains numerous specifications of third-party APIs for each vulnerability class. However, as we show later in our evaluation, despite having such specialized queries and extensive specifications, CodeQL fails to detect a majority of vulnerabilities in real-world projects. For our analysis, we write a specialized CodeQL query for each vulnerability that uses our mined specifications instead of those provided by CodeQL. Our query for Path Traversal vulnerability (CWE 22) is shown in Listing 3 in the appendix. We develop similar queries for each CWE that we evaluate.
@@ 3.5 Triaging of Alerts via Contextual Analysis
Fig. 4 presents an example prompt for contextual analysis. The prompt includes CWE information and code snippets for nodes along the path, with an emphasis on the source and sink. Specifically, we include ±$\pm 5$ lines surrounding the exact source and sink location, as well as the enclosing function and class. The exact line of source and sink is marked with a comment. For the intermediate steps, we include the file names and the line of code. When the path is too long, we keep only a subset of nodes to limit the size of the prompt. As such, we provide the full context for the potential vulnerability to be thoroughly analyzed.
We expect the LLM to respond in JSON format with the final verdict as well as an explanation to the verdict. The JSON format prompts the LLM to generate the explanation before delivering the final verdict, as presenting the judgment after the reasoning process is known to yield better results. In addition, if the verdict is false, we ask the LLM to indicate whether the source or sink is a false positive, which helps to prune other paths and thereby save on the number of calls to the LLM.
Figure 4: LLM user prompt and response for contextual analysis of data-flow paths. In the user prompt, we mark with color the CWE and path information that is filling the prompt template. For cleaner presentation, we modify the snippets and left out the system prompt.
@@ 3.6 Evaluation Metrics
#$\begin{array}[]{rlrl}\footnotesize\textit{\#VulPath}(P)=&|\{\textit{Path}\in\textit{Paths}^{P}~{}|~{}\textit{Path}~{}\cap~{}\mathbf{V}_{\text{vul}}^{P}\neq\emptyset\}|,&\textit{Rec}(P)=&\mathbbm{1}_{\textit{\#VulPath}(P)>0},\\ \textit{\#Detected}(\mathcal{D})=&\textstyle\sum_{P\in\mathcal{D}}\textit{Rec}(P),&\textit{Prec}(P)=&\textstyle\frac{\textit{\#VulPath}(P)}{|\textit{Paths}^{P}|},\\ \textit{AvgFDR}(\mathcal{D})=&\text{Avg}_{P\in\mathcal{D},|\textit{Paths}^{P}|>0}1-\textit{Prec}(P),&\textit{AvgF1}(\mathcal{D})=&\textstyle\frac{1}{|\mathcal{D}|}\textstyle\sum_{P\in\mathcal{D}}\frac{2\cdot\textit{Prec}(P)\cdot\textit{Rec}(P)}{\textit{Prec}(P)+\textit{Rec}(P)}\end{array}$
We note that $\textit{Prec}(P)$ might sometimes be undefined due to division-by-zero if the detection tool retrieves no path ($|\textit{Paths}^{P}|=0$). Therefore, for AvgFDR to be meaningful, we only consider the projects where the detection tool produces at least one positive result. On the other hand, AvgF1 would not suffer from this problem because $\textit{Rec}(P)=0$ when no positive result is returned, making the whole F1 term 00 regardless of undefinedness of $\textit{Prec}(P)$.
@@ 4 CWE-Bench-Java: A Dataset of Security Vulnerabilities in Java
To evaluate our approach, we require a dataset of vulnerable versions of Java projects with several important characteristics: 1) Each benchmark should have relevant vulnerability metadata, such as the CWE ID, CVE ID, fix commit, and vulnerable project version, 2) each project in the dataset must be compilable, which is a key requirement for static analysis and data flow graph extraction, 3) the projects must be real-world, which are typically more complex and hence challenging to analyze compared to synthetic benchmarks, and 4) finally, each vulnerability and its location (e.g., method) in the project must be validated so that this information can be used for robust evaluation of vulnerability detection tools. Unfortunately, no existing dataset satisfies all these requirements. Table 5 presents a comparison of our dataset, which we discuss next, with prior vulnerability datasets.
@@
To address these requirements, we curate our own dataset of vulnerabilities. For this paper, we focus only on vulnerabilities in Java libraries that are available via the widely used Maven package manager. We choose Java because it is commonly used to develop server-side, Android, and web applications, which are prone to security risks. Further, due to Java’s long history, there are many existing CVEs in numerous Java projects that are available for analysis. We initially use the GitHub Advisory database [15, 16] to obtain such vulnerabilities, and further filter it with cross-validated information from multiple sources, including manual verification. Fig. 5 illustrates the complete set of steps for curating CWE-Bench-Java.
@@ 5 Evaluation
We perform extensive experimental evaluations of IRIS and demonstrate its practical effectiveness in detecting vulnerabilities in real-world Java repositories in CWE-Bench-Java. Due to space limits, we include additional results and analyses in the appendix. We answer the following research questions:
@@ 5.1 Experimental Setup
We select two closed-source LLMs from OpenAI: GPT 4 (gpt-4-0125-preview) and GPT 3.5 (gpt-3.5-turbo-0125) for our evaluation. We also select instruction-tuned versions of three open-source LLMs via huggingface API: Llama 3 8B and 70B, and DeepSeekCoder 7B. For the CodeQL baseline, we use version 2.15.3 and its built-in Security queries specifically designed for each CWE. Other baselines included are Facebook Infer [11], SpotBugs [26], and Snyk [39]. We expand further on the other experimental setups in the appendix.
@@ 5.2 RQ1: Effectiveness of IRIS on Detecting Existing Vulnerabilities
Table 1: Overall performance comparison of CodeQL vs IRIS on Detection Rate ($\uparrow$), Average FDR ($\downarrow$), and Average F1 ($\uparrow$). We present results of IRIS with different LLMs including OpenAI GPT-4 and GPT-3.5, Llama-3 (L3) 8B and 70B, and DeepSeekCoder (DSC) 7B.
@@
Effectiveness of IRIS. The results in Table 1 highlight IRIS’s superior performance compared to CodeQL. Specifically, IRIS, when paired with GPT-4, identifies 55 vulnerabilities—28 more than CodeQL. While GPT-4 shows the highest efficacy, smaller, specialized LLMs like DeepSeekCoder 7B still detect 52 vulnerabilities, suggesting that our approach can effectively leverage smaller-scale models, enhancing accessibility. Notably, this increase in detected vulnerabilities does not compromise precision, as evidenced by IRIS’s lower average false discovery rate (FDR) with GPT-4 compared to CodeQL. Moreover, IRIS improves average F1 by 0.1, reflecting a better balance between precision and recall. We note that the reported average FDR is an upper bound, as our metrics may overlook other true vulnerabilities in the repository. To further assess detection accuracy, we randomly sampled 50 alarms reported by IRIS using GPT-4, and found that 27 out of 50 exhibit potential attack surfaces, yielding a more refined estimated false discovery rate of 46%.
Table 2 presents a detailed breakdown of detected vulnerabilities, comparing IRIS against various baselines. With the exception of IRIS using Llama-3 8B, which underperforms in detecting CWE-22 vulnerabilities, IRIS consistently outperforms all other baselines. Notably, CWE-78 (OS Command Injection) remains particularly challenging for all LLMs. Our manual investigation revealed that the vulnerability patterns in CWE-78 are highly intricate, often involving OS command injections via gadget-chains [2] or external side effects, such as file writes, which are difficult to track using static analysis. This highlights the inherent limitations of static analysis, as opposed to dynamic approaches—an area that we leave for future work.

### v3

# IRIS: LLM-Assisted Static Analysis for Detecting Security Vulnerabilities

@@ Abstract
Software is prone to security vulnerabilities. Program analysis tools to detect them have limited effectiveness in practice due to their reliance on human labeled specifications. Large language models (or LLMs) have shown impressive code generation capabilities but they cannot do complex reasoning over code to detect such vulnerabilities especially since this task requires whole-repository analysis. We propose IRIS, a neuro-symbolic approach that systematically combines LLMs with static analysis to perform whole-repository reasoning for security vulnerability detection. Specifically, IRIS leverages LLMs to infer taint specifications and perform contextual analysis, alleviating needs for human specifications and inspection. For evaluation, we curate a new dataset, CWE-Bench-Java, comprising 120 manually validated security vulnerabilities in real-world Java projects. A state-of-the-art static analysis tool CodeQL detects only 27 of these vulnerabilities whereas IRIS with GPT-4 detects 55 ($+28$) and improves upon CodeQL’s average false discovery rate by 5% points. Furthermore, IRIS identifies 4 previously unknown vulnerabilities which cannot be found by existing tools. IRIS is available publicly at https://github.com/iris-sast/iris.
@@ 1 Introduction
Security vulnerabilities pose a major threat to the safety of software applications and its users. In 2023 alone, more than 29,000 CVEs were reported—almost 4000 higher than in 2022 (CVE Trends, ). Detecting vulnerabilities is extremely challenging despite advances in techniques to uncover them. A promising such technique called static taint analysis is widely used in popular tools such as GitHub CodeQL (Avgustinov et al., 2016), Facebook Infer (FB Infer, ), Checker Framework (Checker Framework, ), and Snyk Code (Snyk.io, ). These tools, however, face several challenges that greatly limit their effectiveness and accessibility in practice.
@@
False positives due to lack of precise context-sensitive and intuitive reasoning. Second, it is well-known that static analysis often suffers from low precision, i.e., it may generate many false alarms (Kang et al., 2022; Johnson et al., 2013). Such imprecision stems from multiple sources. For instance, the source or sink specifications may be spurious, or the analysis may over-approximate over branches in code or possible inputs. Further, even if the specifications are correct, the context in which the detected source or sink is used may not be exploitable. Hence, a developer may need to triage through several potentially false security alerts, wasting significant time and effort.
Limitations of prior data-driven approaches to improve static taint analysis. Many techniques have been proposed to address the challenges of static taint analysis. For instance, Livshits et al. (2009) proposed a probabilistic approach, MERLIN, to automatically mine taint specifications. A more recent work, Seldon (Chibotaru et al., 2019), improves the scalability of this approach by formulating the taint specification inference problem as a linear optimization task. However, such approaches rely on analyzing the code of third-party libraries to extract specifications, which is expensive and hard to scale. Researchers have also developed statistical and learning-based techniques to mitigate false positive alerts (Jung et al., 2005; Heckman & Williams, 2009; Hanam et al., 2014). However, such approaches still have limited effectiveness in practice (Kang et al., 2022).
Large Language Models (or LLMs) have made impressive strides in code generation and summarization. LLMs have also been applied to code related tasks such as program repair (Xia et al., 2023), code translation (Pan et al., 2024), test generation (Lemieux et al., 2023), and static analysis (Li et al., 2024). Recent studies (Steenhoek et al., 2024; Khare et al., 2023) evaluated LLMs’ effectiveness at detecting vulnerabilities at the method level and showed that LLMs fail to do complex reasoning with code, especially because it depends on the context in which the method is used in the project. On the other hand, recent benchmarks like SWE-Bench (Jimenez et al., 2023) show that LLMs are also poor at doing project-level reasoning. Hence, an intriguing question is whether LLMs can be combined with static analysis to improve their reasoning capabilities. In this work, we answer this question in the context of vulnerability detection and make the following contributions:
@@
Dataset. We curate a dataset of manually vetted and compilable Java projects, CWE-Bench-Java, containing 120 vulnerabilities (one per project) across four common vulnerability classes. The projects in the dataset are complex, containing 300K lines of code on average, and 10 projects with more than a million lines of code each, making it a challenging benchmark for vulnerability detection. The dataset and the corresponding scripts to fetch, build, and analyze the Java projects are available publicly at https://github.com/iris-sast/cwe-bench-java.
Results. We evaluate IRIS on CWE-Bench-Java using 7 diverse open- and closed-source LLMs. Overall, IRIS obtains the best results with GPT-4, detecting 55 vulnerabilities, which is 28 ($103.7\%$) more than CodeQL, the existing best-performing static analyzer. We show that the increase is not at the expense of false positives, as IRIS with GPT-4 achieves an average false discovery rate of $84.82\%$, which is $5.21\%$ lower than that of CodeQL. Further, when applied to the latest versions of 30 Java projects, IRIS with GPT-4 discovered 4 previously unknown vulnerabilities.
@@ 2 Motivating Example
In contrast, IRIS takes a different approach by inferring project- and vulnerability-specific specifications on-the-fly by using LLMs. The LLM-based components in IRIS correctly identify the untrusted source and the vulnerable sink. IRIS augments CodeQL with these specifications and successfully detects the unsanitized dataflow path between the detected source and sink in the repository. However, augmented CodeQL produces many false positives, which are hard to eliminate using logical rules. To solve this challenge, IRIS encodes the detected code paths and the surrounding context into a simple prompt and uses an LLM to classify it as true or false positive. Specifically, out of 8 paths reported by static analysis, 5 false positives are filtered out, leaving the path in Fig. 2 as one of the final alarms. Overall, we observe that IRIS can detect many such vulnerabilities that are beyond the reach of CodeQL-like static analysis tools, while keeping false alarms to a minimum.
@@ 3.3 Inferring Taint Specifications using LLMs
Due to the sheer number of APIs to be labeled, we insert a batch of APIs in a single prompt and ask the LLM to respond with JSON formatted strings. The batch size is a tunable hyper-parameter. We adopt few-shot (usually 3-shot) prompting strategy for labeling external APIs, while zero-shot is used for labeling internal APIs. Notably for internal APIs, we also include information from repository readme and JavaDoc documentations, if applicable. In practice, we find that this extra information helps LLM understand the high-level purpose and usage of the codebase, resulting in better labeling accuracy. At the end of this stage, we have successfully obtained $\bm{S}_{\textit{source}}^{C}$ and $\bm{S}_{\textit{sink}}^{C}$ which are going to be used by the static analysis engine in the next stage.
Figure 4: LLM user prompt and response for contextual analysis of dataflow paths. In the user prompt, we mark with color the CWE and path information that is filling the prompt template. For cleaner presentation, we modify the snippets and left out the system prompt.
@@ 3.4 Vulnerability Detection
$\textit{CodeQL}(\mathbb{G}^{P},\bm{S}_{\textit{source}}^{C},\bm{S}_{\textit{sink}}^{C},\textit{Query}^{C})=\{\textit{Path}_{1},\ldots,\textit{Path}_{k}\}.$
CodeQL itself contains numerous specifications of third-party APIs for each vulnerability class. However, as we show later in our evaluation, despite having such specialized queries and extensive specifications, CodeQL fails to detect a majority of vulnerabilities in real-world projects. For our analysis, we write a specialized CodeQL query for each vulnerability that uses our mined specifications instead of those provided by CodeQL. Details of our queries are described in the Appendix.
@@ 3.5 Triaging of Alerts via Contextual Analysis
Fig. 4 presents an example prompt for contextual analysis. The prompt includes CWE information and code snippets for nodes along the path, with an emphasis on the source and sink. For the intermediate steps, we include the file names and the line of code. When the path is too long, we keep only a subset of nodes to limit the size of the prompt. More details and design decisions of this prompt are described in the Appendix. We expect the LLM to respond in JSON format with the final verdict as well as an explanation to the verdict. The JSON format prompts the LLM to generate the explanation before delivering the final verdict, as presenting the judgment after the reasoning process is known to yield better results. In addition, if the verdict is false, we ask the LLM to indicate whether the source or sink is a false positive, which helps to prune other paths and thereby save on the number of calls to the LLM.
@@ 3.6 Evaluation Metrics
#$\begin{array}[]{rlrl}\textit{\#VulPath}(P)=&|\{\textit{Path}\in\textit{Paths}^{P}~{}|~{}\textit{Path}~{}\cap~{}\mathbf{V}_{\text{vul}}^{P}\neq\emptyset\}|,&\textit{Rec}(P)=&\mathbbm{1}_{\textit{\#VulPath}(P)>0},\\ \textit{\#Detected}(\mathcal{D})=&\textstyle\sum_{P\in\mathcal{D}}\textit{Rec}(P),&\textit{Prec}(P)=&\textstyle\frac{\textit{\#VulPath}(P)}{|\textit{Paths}^{P}|},\\ \textit{AvgFDR}(\mathcal{D})=&\text{avg}_{P\in\mathcal{D},|\textit{Paths}^{P}|>0}1-\textit{Prec}(P),&\textit{AvgF1}(\mathcal{D})=&\textstyle\frac{1}{|\mathcal{D}|}\textstyle\sum_{P\in\mathcal{D}}\frac{2\cdot\textit{Prec}(P)\cdot\textit{Rec}(P)}{\textit{Prec}(P)+\textit{Rec}(P)}.\end{array}$
Specifically, a lower AvgFDR is preferable, as it indicates a lower ratio of false positives. We note that $\textit{Prec}(P)$ might sometimes be undefined due to division-by-zero if the detection tool retrieves no path ($|\textit{Paths}^{P}|=0$). Therefore, for AvgFDR to be meaningful, we only consider the projects where at least one positive result is produced ($|\textit{Paths}^{P}|>0$). AvgF1 avoids this issue since $\textit{Rec}(P)=0$ when no positive labels exist, forcing the F1 term to be zero regardless of $\textit{Prec}(P)$.
@@ 4 CWE-Bench-Java: A Dataset of Security Vulnerabilities in Java
To evaluate our approach, we require a dataset of vulnerable versions of Java projects with several important characteristics: 1) Each benchmark should have relevant vulnerability metadata, such as the CWE ID, CVE ID, fix commit, and vulnerable project version, 2) each project in the dataset must be compilable, which is a key requirement for static analysis and data flow graph extraction, 3) the projects must be real-world, which are typically more complex and hence challenging to analyze compared to synthetic benchmarks, and 4) finally, each vulnerability and its location (e.g., method) in the project must be validated so that this information can be used for robust evaluation of vulnerability detection tools. Unfortunately, no existing dataset satisfies all these requirements.
@@
To address these requirements, we curate our own dataset of vulnerabilities. For this paper, we focus only on vulnerabilities in Java libraries that are available via the widely used Maven package manager. We choose Java because it is commonly used to develop server-side, Android, and web applications, which are prone to security risks. Further, due to Java’s long history, there are many existing CVEs in numerous Java projects that are available for analysis. We initially use the GitHub Advisory database (GitHub, 2024b; c) to obtain such vulnerabilities, and further filter it with cross-validated information from multiple sources, including manual verification. Fig. 5 illustrates the complete set of steps for curating CWE-Bench-Java.
@@ 5 Evaluation
We perform extensive experimental evaluations of IRIS and demonstrate its practical effectiveness in detecting vulnerabilities in real-world Java repositories in CWE-Bench-Java. While additional results and analyses are provided in the Appendix, we address the following key research questions:
@@ 5.1 Experimental Setup
We select two closed-source LLMs from OpenAI: GPT-4 (gpt-4-0125-preview) and GPT-3.5 (gpt-3.5-turbo-0125) for our evaluation. We also select instruction-tuned versions of four open-source LLMs via huggingface API: Llama 3 (L3) 8B and 70B, Qwen-2.5-Coder (Q2.5C) 32B, Gemma-2 (G2) 27B, and DeepSeekCoder (DSC) 7B. For the CodeQL baseline, we use version 2.15.3 and its built-in Security queries specifically designed for each CWE. Other baselines included are Facebook Infer (FB Infer, ), SpotBugs (Lavazza et al., 2020), and Snyk (Snyk.io, ). We expand further on the other experimental setups in the Appendix.
@@ 5.2 RQ1: Effectiveness of IRIS on Detecting Existing Vulnerabilities
Table 1: Overall performance comparison of CodeQL vs IRIS on Detection Rate ($\uparrow$), Average FDR ($\downarrow$), and Average F1 ($\uparrow$). We present results of IRIS with LLMs including GPT-4 and GPT-3.5, L3 8B and 70B, Q2.5C 32B, G2 27B, and DSC 7B.
@@
Q2.5C 32B $47$ ($\uparrow 20$) $39.17$ ($\uparrow 16.67$) $92.38$ ($\uparrow 2.35$) $0.097$ ($\uparrow 0.021$)
G2 27B $45$ ($\uparrow 18$) $37.50$ ($\uparrow 15.00$) $91.23$ ($\uparrow 1.20$) $0.100$ ($\uparrow 0.024$)
@@
The results in Table 1 highlight IRIS’s superior performance compared to CodeQL. Specifically, IRIS, when paired with GPT-4, identifies 55 vulnerabilities—28 more than CodeQL. While GPT-4 shows the highest efficacy, smaller, specialized LLMs like DeepSeekCoder 7B still detect 52 vulnerabilities, suggesting that our approach can effectively leverage smaller-scale models, enhancing accessibility. Notably, this increase in detected vulnerabilities does not compromise precision, as evidenced by IRIS’s lower average false discovery rate (FDR) with GPT-4 compared to CodeQL. Moreover, IRIS improves average F1 by 0.1, reflecting a better balance between precision and recall. We note that the reported average FDR is a coarse measure as our metrics may consider a true (but unknown) vulnerability found by IRIS as a false positive. Hence, the reported FDR is an upper bound. To get a better sense of detection accuracy, we manually analyzed 50 random alarms reported by IRIS (using GPT-4) and found that 27/50 alarms exhibit potential attack surfaces, yielding a more refined estimated false discovery rate of 46%. Hence, IRIS will likely be more effective in practice.
Table 2 presents a detailed breakdown of detected vulnerabilities, comparing IRIS against various baselines. With the exception of IRIS using Llama-3 8B, which underperforms in detecting CWE-22 vulnerabilities, IRIS consistently outperforms all other baselines. Notably, CWE-78 (OS Command Injection) remains particularly challenging for all LLMs. Our manual investigation revealed that the vulnerability patterns in CWE-78 are highly intricate, often involving OS command injections via gadget-chains (Cao et al., 2023) or external side effects, such as file writes, which are difficult to track using static analysis. This highlights the inherent limitations of static analysis, as opposed to dynamic approaches—an area that we leave for future work.